# Joint-Embedding Masked Autoencoder for Self-supervised Learning of Dynamic Functional Connectivity from the Human Brain

### v1

@@ VI-D Ablation Study
VI-D1 Fine-tuning Performance with Limited Samples
In learning algorithms for analyzing or diagnosing brain disorders with clinical fMRI datasets, a common challenge is the limited number of available training samples. To address this, we aim to enhance model performance in scenarios with limited samples through self-supervised learning. We assessed this by randomly sampling $10\%$, $20\%$, $30\%$, and $50\%$ of the total 884 samples from the ABIDE dataset, and compared the performance with other dynamic graph SSL methods on psychiatric diagnosis classification. As shown in Fig. 2, our proposed ST-JEMA demonstrated superior classification performance even with fewer samples compared to other approaches. Specifically, with the utilization of 20% and 30%, ST-JEMA significantly outperforms other baseline methods. This indicates that the node representations pre-trained with ST-JEMA are more useful for performing downstream tasks, evidencing its potential as a solid pre-trained model for developing more data-efficient models. Notably, ST-DGI fails to outperform Random-Init when utilizing less than 50% of the downstream dataset. This suggests that the pre-trained representation space established through contrastive-based SSL encounters challenges in adapting to the downstream task with a limited number of samples.
VI-D2 Robustness on Missing Data Scenarios
VI-D3 Ablation on Loss Function
@@
TABLE VI: Ablation results on reconstruction strategies in the loss ${\mathcal{L}}_{\text{ST-JEMA}}$. Here, Baseline (1), (2), (3), and (4) refer to the scenarios where node representation reconstruction, edge reconstruction, spatial, and temporal loss are excluded from ST-JEMA, respectively.
@@
Method Node Edge ${\mathcal{L}}_{spatial}$ ${\mathcal{L}}_{temporal}$ AUROC ($\uparrow$) Random Init 67.16 Baseline (1) 68.24 Baseline (2) 70.33 Baseline (3) 69.79 Baseline (4) 65.19 ST-JEMA (Ours) $\mathbf{71.49}$
@@
In formulating the loss ${\mathcal{L}}_{\text{ST-JEMA}}$ in ST-JEMA, four crucial components are considered: 1) node representation reconstruction loss, 2) edge reconstruction loss, 3) spatial loss, and 4) temporal loss. We perform an ablation study on the construction of the loss function to prove the significance of each component in ${\mathcal{L}}_{\text{ST-JEMA}}$. In this ablation experiment, our target task is the psychiatric diagnosis classification task within the ABIDE dataset. Table VI clearly shows that excluding any loss term from ${\mathcal{L}}_{\text{ST-JEMA}}$ results in a significant drop in performance compared to ST-JEMA. Notably, the most substantial performance decline is observed when excluding the temporal loss ${\mathcal{L}}_{\text{temp}}$, highlighting the critical importance of capturing temporal dynamics in fMRI data for accurately addressing downstream tasks. Meanwhile, the performance of baselines, except when excluding the temporal loss, consistently outperforms the Random Init baseline. This demonstrates the significance of pre-training on a large upstream dataset, particularly in scenarios where the downstream data is limited in terms of labeled samples, as commonly encountered in clinical fMRI datasets. Refer to Appendix Appendix E to see further ablation experiments on the impact of pre-training data size, block mask ratio, and decoder architecture.
VII Conclusion
In this paper, we introduced the Spatio-Temporal Joint-Embedding Masked Autoencoder (ST-JEMA), SSL framework designed for fMRI datasets. Leveraging the recent SSL framework for vision tasks known as JEPA [2], ST-JEMA effectively captures both spatial and temporal dynamics in the dynamic FC of fMRI data by concurrently incorporating two types of reconstruction losses: ${\mathcal{L}}_{\text{spat}}$ and ${\mathcal{L}}_{\text{temp}}$. We empirically validate that ST-JEMA enhances the performance on non-clinical and clinical datasets and three types of tasks compared to competing SSL methods. These findings prove the efficacy of ST-JEMA in fMRI downstream tasks, particularly in scenarios where labeled datasets are limited, which is a common situation in the fMRI research community.
@@ Acknowledgments
We would like to thank Sungho Keum for his valuable discussions that helped in advancing our ideas. This research was partly supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (NRF-2022R1I1A1A01069589), National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (NRF-2021M3E5D9025030), Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.2019-0-00075, Artificial Intelligence Graduate School Program(KAIST), No.2022-0-00713, Meta-learning Applicable to Real-world Problems).
@@ Appendix A Algorithmic flow of ST-JEMA
To further facilitate understanding, we summarized the overall pipeline of our proposed Spatio-Temporal Joint Embedding Masked Autoencoder (ST-JEMA) framework for self-supervised learning on fMRI data in Algorithm 1. The algorithm describes the spatial and temporal reconstruction process of node representations and adjacency matrices employing K block masks, assuming the initial conversion of fMRI BOLD signals $\boldsymbol{P}$ into dynamic graphs ${\mathcal{G}}(t)$ has already been completed. Detailed explanations of each equation and the methodological underpinnings can be found in Section IV.
@@
17:         $\tilde{\boldsymbol{Z}}_{\texttt{cxt}}(t)\leftarrow\tilde{\boldsymbol{Z}}\odot\Big{(}\boldsymbol{M}_{\boldsymbol{X}}^{(1)}(t)\odot\boldsymbol{M}_{\boldsymbol{X}}^{(2)}(t)\odot\ldots\odot\boldsymbol{M}_{\boldsymbol{X}}^{(K)}(t)\Big{)}$
@@
Appendix B Datasets
TABLE VII: Phenotype information summary of the non-clinical and clinical rs-fMRI datasets
@@ B-A Data Preprocessing
From the time-series BOLD signal $\boldsymbol{P}$ in each fMRI dataset, we construct two types of graphs: a static graph and a dynamic graph. While the dynamic graph is constructed as described in Section III-C, for the static graph, we build the adjacency matrix $\boldsymbol{A}$ to represent the functional connectivity of ROI by windowing across all time stamps $\tau\in[T_{max}]$. Additionally, the node feature $\boldsymbol{X}$ is constructed using one-hot positional encoding, followed by a projection through the matrix $\boldsymbol{W}$.
B-B Data Description
UK Biobank (UKB) [23]: We utilized 40,913 samples of resting-state fMRI from the UK Biobank, a comprehensive biomedical database of genetic, lifestyle, and health information from half a million participants. This dataset supports advanced medical research by offering insights into serious illnesses through diverse data collections, including genome sequencing, health record linkages, and functional MRI imaging.
Adolescent Brain Cognitive Development (ABCD) [4]: A longitudinal study tracking nearly 12,000 youth across 21 research sites to investigate the impact of childhood experiences on brain development and various behavioral, social, and health outcomes. We constructed 9,111 samples aged 9 to 11 years.
Human Connectome Project Young Adults (HCP-YA) [26]: Mapping the connectome of healthy individuals, this project provides a comprehensive database from 1,200 young adults, incorporating advanced 3T and 7T MRI scanners for an exploration of brain connectivity and function. In our study, we specifically utilized rs-fMRI data from 1,093 subjects, following the pre-processing methodology described in Kim et al. [12].
@@ B-B Data Description
Human Connectome Project Development (HCP-D) [22]: This dataset includes over 1,300 participants ranging from 5 to 21 years old, exploring the dynamics of brain development from childhood through adolescence using a 3T MRI scanner. Specifically, our study utilizes rs-fMRI data from 632 individuals aged between 8 and 21 years.
@@
Autism Brain Imaging Data Exchange (ABIDE) [6]: This dataset gathers rs-fMRI, anatomical, and phenotypic data from 1,112 individuals across 17 international sites, aiming to enhance the understanding of Autism Spectrum Disorder (ASD) through collaborative research. For our study, we specifically utilized rs-fMRI data from 884 individuals, ranging in age from 6 to 64 years.
@@
Appendix C Competing Methods
@@
Random-Init: Training from scratch on the downstream dataset without the pre-training step using the upstream dataset. We present two sets of results for Random-Init, specifically when using a static graph as input and a dynamic graph as input, as elaborated in Section B-A.
@@
Appendix D Experimental Details
@@ D-A Architecture Details
D-B Hyperparameter Details
In this study, we optimize GNN encoder in two phases: 1) pretraining the GNN encoder using SSL method on the upstream dataset, 2) finetuning the GNN encoder for classification or regression tasks on the downstream dataset.
@@ D-B Hyperparameter Details
Additional hyperparameters for each SSL method are summarized as follows:
  • SimGRACE: Perturbation scale $\eta=1.0$.
  • GraphMAE: Random masking ratio $\alpha\,(\%)=\{30,50,70\}$.
  • ST-DGI: $k=\{1,3,6\}$.
  • ST-MAE: Random masking ratio $\alpha\,(\%)=\{30,50,70\}$.
  • ST-JEMA: Block masking ratio $[\alpha_{\min},\alpha_{\max}]\,(\%)=\{10\text{-}15,15\text{-}20,20\text{-}25,25\text{-}30\}$, $\gamma=0.5$, $\lambda_{\text{node}}=1.0$, $\lambda_{\text{edge}}=\{1\times 10^{-3},1\times 10^{-4},1\times 10^{-5},1\times 10^{-6}\}$.
Appendix E Additional Ablation Studies
We conduct additional ablation studies to evaluate the impact of SSL data size, the optimal block mask ratio., and assess the feasibility of the proposed decoder architecture.
E-A Impact of Pre-Training Data Size
@@
E-B Ablation on Block Mask Ratio
@@
E-C Decoder Architecture
TABLE VIII: Ablation results on different architecture for decoder module with diagnosis classification task on ABIDE dataset. The performance is evaluated using the AUROC score, with higher scores indicating better performance.
@@ E-C Decoder Architecture
As described in Section IV, We implemented the MLP-Mixer [24] to reconstruct the target node representation by integrating information from context node representations. Another approach to aggregate representations from different nodes is by using a GNN decoder. GraphMAE employed a more expressive GNN decoder instead of a simple MLP to prevent learning representations nearly identical to the low-level semantics of input features, thus aiming to capture higher-level semantics.
To determine which decoder structure is most effective for our methodology, we compared an MLP module, a one-layer GIN module, and a one-layer MLP-Mixer module. The MLP module does not utilize information from other node representations for reconstructing the target node and hence does not use a learnable token for node representation; instead, it reconstructs the target node representation using the corresponding node representation obtained from the GNN encoder. As seen in Table VIII, the MLP decoder exhibited the lowest performance, possibly due to the pre-trained representations being of low-level semantics. While the GIN module performed better, its performance was similar to that of a random initialization, indicating limited benefit. In contrast, our method using the MLP-Mixer facilitated interactions among context node representations, thereby learning representations with higher-level semantics.
Appendix F Additional Related Works
F-A Contrastive Self-supervised Learning on Graph
@@
F-B Generative Self-supervised Learning on Graph

### v2

@@ 6.2 Ablation Study
6.2.1 Fine-tuning Performance with Limited Samples
Clinical fMRI datasets often suffer from limited labeled data, which hinders the training of robust models. To assess the effectiveness of our pre-trained GNN under such constraints, we conducted experiments on $10\%$, $20\%$, $30\%$, and $50\%$ subsets of the ABIDE dataset. and compared ST-JEMA with other dynamic graph SSL methods on psychiatric diagnosis classification. As shown in Fig. 2, ST-JEMA consistently outperformed all baselines, even with as little as $20\%$ of the training data, indicating its ability to learn semantic representations beneficial under data scarcity. Notably, ST-DGI failed to surpass Random-Init when using less than $50\%$, revealing the limitations of contrastive-based methods. These results demonstrate the effectiveness of generative SSL for clinical neuroimaging tasks where labeled data is scarce.
6.2.2 Ablation on Loss Function
@@
Table 4: Ablation results on reconstruction losses in ${\mathcal{L}}_{\text{ST-JEMA}}$. Baseline (1)–(4) node, edge, spatial, and temporal loss components, respectively.
@@
Method Node Edge ${\mathcal{L}}_{spatial}$ ${\mathcal{L}}_{temporal}$ AUROC ($\uparrow$)
Random Init 67.16
Baseline (1) 68.24
Baseline (2) 70.33
Baseline (3) 69.79
Baseline (4) 65.19
ST-JEMA (Ours) $\mathbf{71.49}$
@@
In formulating the loss ${\mathcal{L}}_{\text{ST-JEMA}}$, we consider four key components: 1) node representation reconstruction, 2) edge reconstruction, 3) spatial, and 4) temporal objectives. We evaluate their contributions through an ablation study on the psychiatric diagnosis classification task using the ABIDE dataset. Table 4 clearly shows that excluding any loss term from ${\mathcal{L}}_{\text{ST-JEMA}}$ results in a significant drop in performance compared to ST-JEMA. Notably, the most substantial performance decline is observed when excluding the temporal loss ${\mathcal{L}}_{\text{temp}}$, highlighting the critical importance of capturing temporal dynamics in fMRI data for accurately addressing downstream tasks. Meanwhile, the performance of baselines, except when excluding the temporal loss, consistently outperforms the Random Init baseline. This demonstrates the significance of pre-training on a large upstream dataset, particularly in scenarios where the downstream data is limited in terms of labeled samples, as commonly encountered in clinical fMRI datasets.
Refer to Appendix F to see further ablation experiments on the impact of pre-training data size, block mask ratio, and decoder architecture.
7 Conclusion
We introduced Spatio-Temporal Joint-Embedding Masked Autoencoder (ST-JEMA), SSL framework tailored for fMRI data. Leveraging the JEPA [14] framework, ST-JEMA captures both spatial and temporal dynamics in the dynamic FC of fMRI data by jointly optimizing spatial and temporal reconstruction objectives. Extensive experiments on both non-clinical and clinical downstream tasks demonstrate that our method outperforms existing SSL approaches across classification and regression tasks. These findings highlight the effectiveness of ST-JEMA in learning robust representations from unlabeled fMRI data, particularly in data-constrained scenarios common in neuroimaging research.
@@ Acknowledgments
We would like to thank Sungho Keum for his valuable discussions that contributed to the development of this work. This research was partly supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (NRF-2022R1I1A1A01069589), National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (NRF-2021M3E5D9025030), Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.2019-0-00075, Artificial Intelligence Graduate School Program(KAIST), No.2022-0-00713, Meta-learning Applicable to Real-world Problems).
@@ Appendix A Algorithmic Flow of ST-JEMA
To further facilitate understanding, we summarized the overall pipeline of our proposed Spatio-Temporal Joint Embedding Masked Autoencoder (ST-JEMA) framework for self-supervised learning on fMRI data in Algorithm 1. The algorithm describes the spatial and temporal reconstruction process of node representations and adjacency matrices employing K block masks, assuming the initial conversion of fMRI BOLD signals $\boldsymbol{P}$ into dynamic graphs ${\mathcal{G}}(t)$ has already been completed. Detailed explanations of each equation and the methodological underpinnings can be found in Section 4.
@@
17:         $\tilde{\boldsymbol{Z}}_{\texttt{cxt}}(t)\leftarrow\tilde{\boldsymbol{Z}}\odot\bigcap_{k=1}^{K}\boldsymbol{M}^{(k)}_{\boldsymbol{X}}(t)$
@@
Appendix B Constructing Dynamic Graphs from fMRI data
Figure 3: Overview of dynamic graph construction from fMRI data.
[…]
Appendix C Datasets
Table 5: Phenotype information summary of the non-clinical and clinical rs-fMRI datasets
@@ C.1 Data Preprocessing
From the time-series BOLD signal $\boldsymbol{P}$ in each fMRI dataset, we construct two types of graphs: a static graph and a dynamic graph. While the dynamic graph is constructed as described in Appendix B, for the static graph, we build the adjacency matrix $\boldsymbol{A}$ to represent the functional connectivity of ROI by windowing across all time stamps $\tau\in[T_{max}]$. Additionally, the node feature $\boldsymbol{X}$ is constructed using one-hot positional encoding, followed by a projection through the matrix $\boldsymbol{W}$.
C.2 Data Description
UK Biobank (UKB) [28]: We utilized 40 913 samples of resting-state fMRI from the UK Biobank, a comprehensive biomedical database of genetic, lifestyle, and health information from half a million participants. This dataset supports advanced medical research by offering insights into serious illnesses through diverse data collections, including genome sequencing, health record linkages, and functional MRI imaging.
Adolescent Brain Cognitive Development (ABCD) [29]: A longitudinal study tracking nearly 12 000 youth across 21 research sites to investigate the impact of childhood experiences on brain development and various behavioral, social, and health outcomes. We constructed 9,111 samples aged 9 to 11 years.
Human Connectome Project Young Adults (HCP-YA) [30]: Mapping the connectome of healthy individuals, this project provides a comprehensive database from 1200 young adults, incorporating advanced 3T and 7T MRI scanners for an exploration of brain connectivity and function. In our study, we specifically utilized rs-fMRI data from 1,093 subjects, following the pre-processing methodology described in Kim et al. [5].
@@ C.2 Data Description
Human Connectome Project Development (HCP-D) [32]: This dataset includes over 1300 participants ranging from 5 to 21 years old, exploring the dynamics of brain development from childhood through adolescence using a 3T MRI scanner. Specifically, our study utilizes rs-fMRI data from 632 individuals aged between 8 and 21 years.
@@
Autism Brain Imaging Data Exchange (ABIDE) [34]: This dataset gathers rs-fMRI, anatomical, and phenotypic data from 1112 individuals across 17 international sites, aiming to enhance the understanding of Autism Spectrum Disorder (ASD) through collaborative research. For our study, we specifically utilized rs-fMRI data from 884 individuals, ranging in age from 6 to 64 years.
@@
Appendix D Competing Methods
@@
Random-Init: Training from scratch on the downstream dataset without the pre-training step using the upstream dataset. We present two sets of results for Random-Init, specifically when using a static graph as input and a dynamic graph as input, as elaborated in Section C.1.
@@
Appendix E Experimental Details
@@ E.1 Architecture Details
E.2 Experimental Details
Dynamic Graph Construction Settings
@@ E.2 Experimental Details
Training Configuration and Optimization
We optimize the GNN encoder in two stages: (1) self-supervised pretraining on the upstream dataset, and (2) supervised finetuning on downstream tasks.
Pre-training employed a cosine decay learning schedule over 10 000 steps with a batch size of 16. Fine-tuning optimized the encoder with the task-specific head, using orthogonal regularization and a one-cycle scheduler for 30 epochs with a batch size of 32. Dynamic graphs integrated a mean pooling module for final representation processing. Classification tasks used cross-entropy loss, while age regression utilized MSE loss with normalized age data. The training utilized AMD EPYC processors, with NVIDIA RTX A6000 for pre-training and GeForce RTX 3090 for fine-tuning and ablation studies.
@@
Additional hyperparameters for each SSL method are summarized in  Table 6.
Table 6: Additional hyperparameters used for each SSL method.
Method Key Hyperparameters
SimGRACE Perturbation scale $\eta=1.0$
GraphMAE Masking ratio $\alpha=\{30,50,70\}$ (%)
ST-DGI Temporal steps $k=\{1,3,6\}$
ST-MAE Masking ratio $\alpha=\{30,50,70\}$ (%)
[…]
$\lambda_{\text{node}}=1.0$, $\lambda_{\text{edge}}=\{10^{-3},10^{-4},10^{-5},10^{-6}\}$
Appendix F Additional Experiments
Table 7: Linear probing performance of psychiatric diagnosis classification. The performance is evaluated using the AUROC score, with higher scores indicating better performance.
Pre-training AUROC ($\uparrow$)
HCP-EP ABIDE ADHD200 COBRE
ST-DGI 70.66 53.87 53.32 63.73
ST-MAE 70.78 55.09 52.66 63.97
ST-JEMA $\mathbf{72.97}$ $\mathbf{58.30}$ $\mathbf{54.44}$ $\mathbf{66.39}$
Table 8: Multi-task learning performance on ABIDE and ADHD200 datasets. Here, (-) represents a degree of performance drop, and (+) indicates the degree of performance improvement compared to the performance in each single-task learning scenario.
Pre-training ABIDE
Gender Age Diagnosis
AUROC ($\uparrow$) MAE ($\downarrow$) AUROC ($\uparrow$)
ST-DGI 65.00 (-9.87) 4.24 (-0.08) 59.50 (-9.36)
ST-MAE 62.86 (-11.88) 4.11 (-0.01) 58.25 (-11.48)
ST-JEMA $\mathbf{65.60}$ (-9.78) $\mathbf{4.06}$ (+0.02) $\mathbf{66.10}$ (-5.39)
Pre-training ADHD200
Gender Age Diagnosis
AUROC ($\uparrow$) MAE ($\downarrow$) AUROC ($\uparrow$)
ST-DGI 60.06 (-7.98) 1.99 (-0.08) 55.61 (-0.85)
ST-MAE 60.98 (-10.03) $\mathbf{1.89}$ (-0.03) 50.45 (-6.32)
ST-JEMA $\mathbf{68.29}$ (-3.65) $1.91$ (-0.05) $\mathbf{59.84}$ (+0.95)
F.1 Linear Probing Performance
According to the results in Table 7, ST-JEMA exhibited higher AUROC performance than ST-MAE, suggesting that ST-JEMA captures more semantically rich information beneficial for psychiatric diagnosis classification. On the contrary, ST-DGI, a SSL method based on contrastive learning, exhibited lower performance than generative approaches. This suggests that the representations learned by ST-DGI might not be as beneficial for psychiatric diagnosis classification tasks compared to generative-based SSL methods.
F.2 Multi-task Learning Performance
Table 8 clearly shows that ST-JEMA achieved the best performance across all tasks compared to other baselines except for the age regression task in the ADHD200 dataset. Notably, ST-JEMA exhibits the smallest overall performance decline for each task when compared to the performance in the single-task learning scenario, outperforming other baseline models. In particular, regarding the age regression task in the ABIDE dataset and diagnosis classification task in ADHD200, ST-JEMA demonstrates an improvement in performance in the multi-task learning scenario compared to its performance in the single-task learning scenario. These findings highlight that the pre-trained encoder in ST-JEMA has effectively captured mutual information beneficial for addressing diverse downstream tasks.
F.3 Robustness on Missing Data Scenarios
F.4 Impact of Pre-Training Data Size
@@
F.5 Ablation on Block Mask Ratio
@@
F.6 Decoder Architecture
Table 9: Ablation results on different architecture for decoder module with diagnosis classification task on ABIDE dataset. The performance is evaluated using the AUROC score, with higher scores indicating better performance.
@@ F.6 Decoder Architecture
As described in Section 4, We implemented the MLP-Mixer [26] to reconstruct the target node representation by integrating information from context node representations. Another approach to aggregate representations from different nodes is by using a GNN decoder. GraphMAE employed a more expressive GNN decoder instead of a simple MLP to prevent learning representations nearly identical to the low-level semantics of input features, thus aiming to capture higher-level semantics.
To determine which decoder structure is most effective for our methodology, we compared an MLP module, a one-layer GIN module, and a one-layer MLP-Mixer module. The MLP module does not utilize information from other node representations for reconstructing the target node and hence does not use a learnable token for node representation; instead, it reconstructs the target node representation using the corresponding node representation obtained from the GNN encoder. As seen in Table 9, the MLP decoder exhibited the lowest performance, possibly due to the pre-trained representations being of low-level semantics. While the GIN module performed better, its performance was similar to that of a random initialization, indicating limited benefit. In contrast, our method using the MLP-Mixer facilitated interactions among context node representations, thereby learning representations with higher-level semantics.
Appendix G Additional Related Works
G.1 Contrastive Self-supervised Learning on Graph
@@
G.2 Generative Self-supervised Learning on Graph